OpenStudy (anonymous):

HELP PLZ name the quadratic, linear and constant term of f(x)= (x+3)^2

OpenStudy (anonymous):

plzz help its just alg 2

OpenStudy (anonymous):

ok since \[\Large (a+b^2)=a^2+b^2+2ab\] here a=x b=3 \[\Large (x+3)^2=(x)^2+(3)^2+2(x)(3)\] \[\Large =x^2+9+6x\]

OpenStudy (anonymous):

a typo ! correct formula is \[\Large (a+b)^2=a^2+b^2+2ab\]

OpenStudy (anonymous):

@pitachips did you get it?

OpenStudy (anonymous):

yes so its x^2 + 9 +9x

OpenStudy (anonymous):

oops no

OpenStudy (anonymous):

the tem with power of x two is called quadratic the term with power of x 1 is called linear and ofc you know constant!

OpenStudy (anonymous):

no it is \[\Large =x^2+6x+9\]

OpenStudy (anonymous):

x^2+6x+9

OpenStudy (anonymous):

yes thank you so the quadratic term is 6x?

OpenStudy (anonymous):

no Quadratoc tem is x^2 (its power is 2 linear term is 6x (its power is 1) and 9 is constant !!!!!!!

OpenStudy (anonymous):

ohhhhh ok!!!! thank you!!